# Measuring the Temperature of Extremely Hot Shock-Heated Gas in the Major Merger MACS J0717.5++3745 With Relativistic Corrections to the Sunyaev-Zel’dovich Effect

### v1

@@ 2.2 Bolocam
Bolocam was used to observe MACS J0717.5++3745 in two different photometric bands centered on approximately $140$ and $270$ GHz. The data (Planck Team, 2015) are reduced to surface brightness maps using the methods described in Sayers et al. (2013) and Sayers et al. (2019). Here we make use of these surface brightness images in both bands, which both readily cover the field of view of the SPIRE-FTS at approximately uniform depth. The Bolocam map uncertainty is assumed to follow a Gaussian distribution with a standard deviation estimated from extracting the surface brightness from identical locations in 1,000 random noise realizations of the data. The FWHM of the Bolocam PSF, equal to 59″ at 140 GHz and 33″ at 270 GHz, is comparable to that of the SPIRE-FTS over the frequency range considered in our analysis. Therefore, we do not attempt to correct for any PSF effects in our surface brightness measurements. The precise values of $I_{270}$ and $I_{140}$ used throughout this work are shown in Table 2.
2.3 Cluster Member Redshifts
@@
In the Bolocam bands, the kSZe due to the line-of-sight velocity of the gas can produce a non-negligible signal compared to the rSZe signal we aim to measure, as shown in Figure 1. Due to strong degeneracies between the amplitudes of the kSZe and tSZe signals, along with the gas temperature (which specifies the relativistic corrections to those signals), it is not possible to obtain meaningful constraints on all three parameters with the existing spectral coverage. As demonstrated in Sayers et al. (2013) the line-of-sight gas velocities are consistent with the bulk dark matter velocities estimated from spectroscopy of cluster-member galaxies (Ma et al., 2009). We therefore use external data to place a prior on the gas velocity in the region of the cluster observed by the SPIRE-FTS.
@@
Following the technique of Silich et al. (2024), we obtain a map of the dark matter velocity by spatially binning the cluster-member galaxy spectra. To summarize the analysis, we first define a spatial aperture based on the cluster-member galaxy distribution by convolving a cluster-member galaxy count map constructed via the redshift catalog of Ma et al. (2009) with a 22′ peak-normalized top-hat kernel. This new map is then smoothed via convolution with a 33′ FWHM area-normalized Gaussian kernel. We then draw a contour within which each pixel has $\gtrsim 5$ galaxies. To close the contour at large radii, we define a circular aperture with $r=6.5^{\prime}$ ($\simeq 2.5$ Mpc at $z\simeq 0.5$). At any point where the drawn contour would remain open at this circular boundary, we close the contour with the circular aperture. Within this aperture, we generate the spatial bins for the cluster-member galaxy velocity map by applying a publicly-available Python implementation of the weighted Voronoi tessellation algorithm (Diehl & Statler, 2006; Rhea et al., 2020) to our galaxy counts map, requiring $10$ galaxies per spatial bin to ensure statistical significance and consistency between pixels. We then apply an estimator from Beers et al. (1990) to populate the cluster-member galaxy velocity in each bin.
2.4 X-ray Spectroscopy
@@
2.4.1 Chandra
The reduction and analysis of the Chandra data are described in detail by Mantz et. al (in preparation), using ciao version 4.16 and caldb version 4.11.0. In brief, the archival data (Observation 1655, CXC-DS (2002)) were reprocessed using standard techniques to produce level 2 event files, and filtered to remove periods of enhanced background. The source detection code wavdetect was used to find point-like sources, which were masked in later analysis. Images were produced in the 0.6–7.0 keV band, binned to 0.984′′ resolution, and regions for spectral analysis were defined using the contbin algorithm (Sanders, 2016) with a target signal-to-noise of 55 and a geometric constraint parameter of 2. Spectra in these regions were modeled as the sum of a Galactic soft foreground, constrained by ROSAT All-Sky Survey data (Sabol & Snowden, 2019); emission from active galactic nuclei fainter than the wavdetect detection limit; the quiescent particle-induced background (Suzuki et al., 2021); and absorbed thermal emission from the ICM, using the apec code and adopting an equivalent absorbing hydrogen column density of $N_{H}=6.92\times 10^{20}$ cm-2 (HI4PI Collaboration et al., 2016). Full details of the foreground and background modeling, including systematic allowances marginalized over can be found in Mantz et. al (in preparation). All spectra were fit in the 0.6–7.0 keV observer-frame band using xspec (version 12.12.1c) to evaluate the likelihood and the lmc python package to sample the posterior distribution. The analysis assumed a common ICM metallicity among all regions, allowing the temperatures and normalizations to fit independently. The resulting map of best-fitting temperatures is shown in Figure 2.
2.4.2 XMM-Newton
We use the European Photon Imaging Camera (EPIC) to observe this cluster with XMM-Newton. This instrument uses two types of cameras; pn CCDs and MOS CCDs. We use observation $0672420201$, and the data files were reprocessed using the XMM-Newton Science Analysis System v19.1.0. We employed the tasks emchain and epchain to generate calibrated event files from the raw data. For the pn camera, an out-of-time (OOT) event file was also produced using epchain. To ensure data quality, we excluded all events with PATTERN>>12 for the MOS cameras and PATTERN>>4 for the pn camera. Additionally, events adjacent to CCD edges and those associated with bad pixels were removed by applying FLAG = 0 for all cameras. We also excluded data from CCDs identified as being in an anomalous state, following the criteria described by Kuntz & Snowden (2008). Finally, periods of high background were filtered out using the mos-filter and pn-filter tasks.
@@ 2.4.2 XMM-Newton
[…]
2.5 Extracting Fiducial Values from Tessellated Maps
@@
The maps of cluster member galaxy velocities and the electron temperatures from both Chandra and XMM-Newton are composed of irregular-shaped regions defined by the WVT or contbin procedures. To extract the fiducial values of the velocity fields at each of the SPIRE-FTS pointings we use a PSF-weighted average computed from the entire map. The weighting factor is given by $\frac{B}{\sigma^{2}}$ where $B$ is the fraction of the PSF that couples to a given region and $\sigma$ is the measurement uncertainty on the quantity within that region. The SPIRE-FTS SLW bands’ PSF depends on frequency channel, but we make the approximation that it has a Gaussian shape with a FWHM equal to 35″ for all channels. Conversely, for the temperature maps we opt to use a per-pixel bin weighting because of the fact that the X-ray contbins are chosen to have an approximately equal number of total counts and thus an inverse variance weighted approach would be biased towards lower temperatures. The extracted temperature values and their uncertainties are estimated via a Monte Carlo sampler. We produce $10^{4}$ realizations of the temperature maps where each spatial bin is varied as a Gaussian-distributed draw with mean and standard deviation set by the central value and measurement uncertainty. The fiducial temperatures are then taken to be the median and standard deviation of this set of $10^{4}$ realizations and are shown, alongside the fiducial velocity values, in Table 2.
3 SPIRE-FTS Data Reduction
@@
To produce sub-mm spectra that can be compared against models for the rSZe, the raw interferograms reported by the SPIRE-FTS are corrected for detector and optical effects, calibrated, and mapped onto spectral cubes. Systematic instrumental effects are then modeled and removed. The resulting spectral hypercubes are then corrected for foreground astrophysical emission to produce estimates of the diffuse SZe signal.
3.1 Spectral Intensity Estimation and Data Cuts
@@
Figure 3: The Dark Sky observation described in Table 1 as processed through HIPE following the procedure outlined in §3.1 including the standard LR correction. The widths of the lines represent the per-datum error estimate from HIPE. A systematic pattern associated with beam vignetting is evident in the outer ring of detectors; for example, the average standard deviation of the intensity of these outer ring detectors is $6.3$ MJy sr-1 compared to $1.9$ MJy sr-1 for the inner detectors. There remains a low frequency increase up to $\sim 6$ MJy sr-1 that is consistent with a  $4.5$ K blackbody, as we would expect from the beam intercepting cold optics with slowly varying temperatures.
@@
The HIPE Spectrometer Mapping User Reprocessing Script v.14 is used to generate a spectrum of MACS J0717.5++3745 at each of the 112 pointing positions (7 detectors $\times$ 4 observations $\times$ 4 jiggle positions) using the v14.3 SPIRE-FTS calibration tree. A threshold of 4 is used for the baseline correction, and Level 2 deglitching is performed with the mad algorithm. The spectrum2d object is made with beam diameters estimated from SLWC3 and the extended source calibration correction is applied. In Figure 3 we present calibrated spectra produced using the HIPE pipeline with the parameters described above, including the standard LR-mode correction. Notably, the deviation of the outer ring detectors is about $3{-}4\times$ that for the central detectors, and exhibits strongly correlated low-frequency variation over the passband. This is due to a known vignetting effect in the instrument (Benielli et al., 2014) that we find is too severe to model adequately. We therefore opt to perform our analysis on only the seven central detectors where the vignetting is not present and the SZe signal is largest.
@@
Figure 4: The signal after HIPE processing for each of the seven inner SLW bolometers. The black points denote the SPIRE-FTS spectrum that we retrieve from HIPE, the red points indicate the same for the Dark Sky observations, and the green points indicate the cleaned spectrum obtained from subtracting a scaled version of the Dark Sky data. We estimate the noise covariance of these cleaned data from Equation 5, and the square root of the diagonal elements are shown as error bars.
@@
The process for creating the Dark Sky templates from each detector spectrum proceeds as follows. For each detector $i$ and pointing position $j$, an amplitude scaling is computed according to $A_{ij}=\frac{\vec{s}_{ij}\cdot\vec{d}_{i}}{\vec{d_{i}}\cdot\vec{d}_{i}}\vec{d}_{i}$. The corrected spectrum is then computed from $\vec{s}_{{\rm c},ij}=\vec{s}_{ij}-A_{ij}\vec{d}_{i}$. The $\vec{s}$, $A\vec{d}$ and $\vec{s}_{\rm c}$ for each detector are shown in Figure 4. For the jiggle position 0 data, this template subtraction leaves a residual with a standard deviation of $0.6$ MJySr-1, approximately a factor of three better than what was achieved using the standard correction and with no discernible Marchili-like structure. However, this correction works less well for the three jiggle positions that were not present in the Dark Sky observation, which constitute different optical paths through the instrument, and leaves residual Marchili-like structure due to incomplete cancellation. We therefore opt to exclude data taken in jiggle positions $\{1,2,3\}$ from analysis, which increases the statistical uncertainty by a factor of two but removes a likely uncharacteristic systematic error from the measurement.
@@
Here $\sigma_{I}$ is the per spectral channel uncertainty and $\delta I_{\rm T}$ is the partial derivative with respect to $T_{\rm rSZe}$. The inverse-squared of this is plotted as a function of spectral channel, and for several different values of $T_{\rm rSZe}$, in Figure 5. We conclude the optimal spectral region is 550–900 GHz, as the sensitivity to the rSZe signal outside of this range is small and including those data would significantly increase the measurement error. This range also removes the portions of the spectra that may contain significant instrumental systematics.
Figure 5: The constraining power on $T_{\rm rSZe}$ estimated by the inverse square of equation 3. The solid black vertical lines highlight where the low and high frequency cuts from §3.1 are applied, and outside of which the constraining power is small. While the SZe distortion is spectrally smooth (see Fig. 1), the relativistic corrections exhibit $\sim 100$ GHz-wide spectral deviations whose relative amplitudes change over the falling edge of the SZ spectrum. These differential effects provide a lever arm to distinguish gas at different temperatures independently of X-ray information. Hence, the constraining power peaks at  $640$ and  $800$ GHz. This is informative for cutting the low and high frequency ends of the FTS bandpass, as the regions with very low constraining power, e.g. $\nu<550$, will contribute disproportionately more noise to the measurement.
3.2 Spectral Noise Estimation
A covariance matrix is estimated using a time-wise jackknife resampling that is designed to cancel spatially and spectrally stable signals in our data, e.g from astrophysical sources or instrument thermal emission, while retaining variations due to statistical noise and properties of the instrument varying on the scan time scale. To generate this estimate, two hundred of the four hundred spectra for a given pointing are multiplied by negative 1, and the mean is computed with:
@@ 3.2 Spectral Noise Estimation
Here $\vec{s}^{*}_{k\pm}$ denotes the sum over all four hundred spectra for a given pointing where indices with a minus sign are multiplied by negative one and indices with a plus sign are kept the same. The resulting covariance matrix is estimated by computing the outer product of $\vec{s}_{l}$ with itself,
@@
where $\rm T$} denotes the matrix transpose. This process is repeated $2\times 10^{6}$ times (chosen to estimate the covariance matrices to five significant figures) and averaged. We perform these jackknives independently for each detector, jiggle position (only jiggle position 0 is used in the analysis), and observation, producing $28$ unique covariance matrices. Since we expect multiple observations with the same detector to have consistent systematics we average these by detector down to a total of seven. A representative example showing the correlation matrix, which is estimated from the noise covariance matrix, is shown in Figure 6. The characteristic correlation length implied by the noise covariance is approximately 50 GHz, comparable to the nominal LR spectral resolving power of 25 GHz.
@@
3.3 Unwanted Astrophysical Emission

### v2

@@ II.3 Bolocam
Bolocam was used to observe MACS J0717.5++3745 in two different photometric bands centered on approximately $140$ and $270$ GHz. The data are reduced to surface brightness maps using the methods described in Sayers et al. (2013) and Sayers et al. (2019). Here, we make use of these surface brightness images in both bands, which both readily cover the field of view of the SPIRE-FTS with approximately uniform depth. The Bolocam map uncertainty is assumed to follow a Gaussian distribution with a standard deviation estimated from extracting the surface brightness from identical locations in $1000$ random noise realizations of the data. The FWHM of the Bolocam PSF, equal to 59″ at 140 GHz and 33″ at 270 GHz, is comparable to that of the SPIRE-FTS over the frequency range considered in our analysis. Therefore, we do not attempt to correct for any PSF effects in our surface brightness measurements. The precise values of $I_{270}$ and $I_{140}$ used throughout this work are shown in Table 2.
II.4 cluster-member Redshifts
@@
In the Bolocam bands, the kSZe due to the line-of-sight velocity of the gas can produce a nonnegligible signal compared to the rSZe signal we aim to measure, as shown in Figure 1. Due to strong degeneracies between the amplitudes of the kSZe and tSZe signals, along with the gas temperature (which specifies the relativistic corrections to those signals), it is not possible to obtain meaningful constraints on all three parameters with the existing spectral coverage. As demonstrated in Sayers et al. (2013) the line-of-sight gas velocities are consistent with the bulk dark matter velocities estimated from spectroscopy of cluster-member galaxies (Ma et al., 2009). We therefore use external data to place a prior on the gas velocity in the region of the cluster observed by the SPIRE-FTS.
@@
Following the technique of Silich et al. (2024), we obtain a map of the dark matter velocity by spatially binning the cluster-member galaxy spectra. To summarize the analysis, we first define a spatial aperture based on the cluster-member galaxy distribution by convolving a cluster-member galaxy count map constructed via the redshift catalog of Ma et al. (2009) with a 22′ peak-normalized top-hat kernel. This new map is then smoothed via convolution with a 33′ FWHM area-normalized Gaussian kernel. We then draw a contour within which each pixel has $\gtrsim 5$ galaxies. To close the contour at large radii, we define a circular aperture with $r=6.5^{\prime}$ ($\simeq 2.5$ Mpc at $z\simeq 0.5$). At any point where the drawn contour would remain open at this circular boundary, we close the contour with the circular aperture. Within this aperture, we generate the spatial bins for the cluster-member galaxy velocity map by applying a publicly-available Python implementation of the weighted Voronoi tessellation (WVT) algorithm (Diehl and Statler, 2006; Rhea et al., 2020) to our galaxy counts map, requiring $10$ galaxies per spatial bin to ensure statistical significance and consistency between pixels. We then apply an estimator from Beers et al. (1990) to populate the cluster-member galaxy velocity in each bin.
II.5 X-ray Spectroscopy
@@
II.5.1 Chandra
The reduction and analysis of the Chandra data are described in detail by Mantz et al. (2025), using ciao version 4.16 and caldb version 4.11.0. In brief, the archival data (Observation 1655, CXC-DS (2002)) were reprocessed using standard techniques to produce level 2 event files and filtered to remove periods of enhanced background. The source detection code wavdetect was used to find pointlike sources, which were masked in later analysis. Images were produced in the 0.6–7.0 keV band, binned to 0.984′′ resolution, and regions for spectral analysis were defined using the contbin algorithm (Sanders, 2016) with a target signal-to-noise ratio (S/N) of 55 and a geometric constraint parameter of 2. Spectra in these regions were modeled as the sum of a Galactic soft foreground, constrained by ROSAT All-Sky Survey data (Sabol and Snowden, 2019); emission from active galactic nuclei fainter than the wavdetect detection limit; the quiescent particle-induced background (Suzuki et al., 2021); and absorbed thermal emission from the ICM, using the apec code and adopting an equivalent absorbing hydrogen column density of $N_{H}=6.92\times 10^{20}$ cm-2 (HI4PI Collaboration et al., 2016). Full details of the foreground and background modeling, including systematic allowances marginalized over, can be found in Mantz et al. (2025). All spectra were fit in the 0.6–7.0 keV observer-frame band using xspec (version 12.12.1c) to evaluate the likelihood and the lmc python package to sample the posterior distribution. The analysis assumed a common ICM metallicity among all regions, allowing the temperatures and normalizations to fit independently. The resulting map of best-fitting temperatures is shown in Figure 2.
II.5.2 XMM-Newton
We use the European Photon Imaging Camera to observe this cluster with XMM-Newton. This instrument uses two types of cameras: pn CCDs and MOS CCDs. We use observation $0672420201$, and the data files were reprocessed using the XMM-Newton Science Analysis System v19.1.0. We employed the tasks emchain and epchain to generate calibrated event files from the raw data. For the pn camera, an out-of-time event file was also produced using epchain. To ensure data quality, we excluded all events with PATTERN>>12 for the MOS cameras and PATTERN>>4 for the pn camera. Additionally, events adjacent to CCD edges and those associated with bad pixels were removed by applying FLAG = 0 for all cameras. We also excluded data from CCDs identified as being in an anomalous state, following the criteria described by Kuntz and Snowden (2008). Finally, periods of high background were filtered out using the mos-filter and pn-filter tasks.
@@ II.5.2 XMM-Newton
The temperature map shown in Figure 2 was produced following the methodology described in Lovisari et al. (2024). Spatial regions were defined using the WVT binning algorithm (Diehl and Statler, 2006), with an S/N requirement of $\geq$ 50.
II.6 Extracting Fiducial Values from Tessellated Maps
@@
The maps of cluster-member galaxy velocities and the electron temperatures from both Chandra and XMM-Newton are composed of irregular-shaped regions defined by the WVT or contbin procedures. To extract the fiducial values of the velocity fields at each of the SPIRE-FTS pointings we use a PSF-weighted average computed from the entire map. The weighting factor is given by $\frac{B}{\sigma^{2}}$ where $B$ is the fraction of the PSF that couples to a given region and $\sigma$ is the measurement uncertainty on the quantity within that region. The SPIRE-FTS SLW bands’ PSF depends on frequency channel, but we make the approximation that it has a Gaussian shape with a FWHM equal to 35″ for all channels. For comparison, the range of PSF sizes over the frequencies used in this analysis are $31\arcsec$ to $37\arcsec$ with an average of $34\arcsec$. Conversely, for the temperature maps, we opt to use a per-pixel bin weighting because of the fact that the X-ray contbins are chosen to have an approximately equal number of total counts, and thus an inverse variance weighted approach would be biased toward lower temperatures. The extracted temperature values and their uncertainties are estimated via a Monte Carlo sampler. We produce $10^{4}$ realizations of the temperature maps, where each spatial bin is varied as a Gaussian-distributed draw with mean and standard deviation set by the central value and measurement uncertainty. The fiducial temperatures are then taken to be the median and standard deviation of this set of $10^{4}$ realizations and are shown, alongside the fiducial velocity values, in Table 2.
III SPIRE-FTS Data Reduction
@@
To produce submillimeter spectra that can be compared against models for the rSZe, the raw interferograms reported by the SPIRE-FTS are corrected for detector and optical effects, calibrated, and mapped onto spectral cubes. Systematic instrumental effects are then modeled and removed. The resulting spectral hypercubes are then corrected for foreground astrophysical emission to produce estimates of the diffuse SZe signal.
III.1 Spectral Intensity Estimation and Data Cuts
@@
Figure 3: The Dark Sky observation described in Table 1 as processed through HIPE following the procedure outlined in §III.1 including the standard LR correction. The widths of the lines represent the per-datum error estimate from HIPE. All panels have the same $x$ and $y$ axes, which are labeled only for the bottom left panel. A systematic pattern associated with beam vignetting is evident in the outer ring of detectors; for example, the average standard deviation of the intensity of these outer ring detectors is $6.3$ MJy sr-1 compared to $1.9$ MJy sr-1 for the inner detectors. There remains a low-frequency increase up to $\sim 6$ MJy sr-1 that is consistent with a  $4.5$ K blackbody, as we would expect from the beam intercepting cold optics with slowly varying temperatures.
@@
The HIPE Spectrometer Mapping User Reprocessing Script v.14 is used to generate a spectrum of MACS J0717.5++3745 at each of the 112 pointing positions (7 detectors $\times$ 4 observations $\times$ 4 jiggle positions) using the v14.3 SPIRE-FTS calibration tree. A threshold of 4 is used for the baseline correction, and Level 2 deglitching is performed with the mad algorithm. The spectrum2d object is made with beam diameters estimated from SLWC3 and the extended source calibration correction is applied. In Figure 3 we present calibrated spectra produced using the HIPE pipeline with the parameters described above, including the standard LR-mode correction. Notably, the deviation of the outer ring detectors is about $3{-}4\times$ that for the central detectors, and exhibits strongly correlated low-frequency variation over the passband. This is due to a known vignetting effect in the instrument (Benielli et al., 2014) that we find is too severe to model adequately. We therefore opt to perform our analysis on only the seven central detectors where the vignetting is not present, and the SZe signal is largest.
@@
Figure 4: MACS J0717.5++3745 spectra after HIPE processing for each of the seven inner SLW bolometers. The black points denote the SPIRE-FTS spectrum that we retrieve from HIPE, the red points indicate the same for the Dark Sky observations, and the green points indicate the cleaned spectrum obtained from subtracting a scaled version of the Dark Sky data. All panels share the same $x$ and $y$ axes which is depicted only for the lower-left subplot. We estimate the noise covariance of these cleaned data from Equation 5, and the square root of the diagonal elements are shown as error bars.
@@
The process for creating the Dark Sky templates from each detector spectrum proceeds as follows. For each detector $i$ and pointing position $j$, an amplitude scaling is computed according to $A_{ij}=\frac{\vec{s}_{ij}\cdot\vec{d}_{i}}{\vec{d_{i}}\cdot\vec{d}_{i}}\vec{d}_{i}$, where $\vec{d}_{i}$ is a Dark Sky spectrum and $\vec{s}_{i}$ is a spectrum of MACS J0717.5++3745. The corrected spectrum is then computed from $\vec{s}_{{\rm c},ij}=\vec{s}_{ij}-A_{ij}\vec{d}_{i}$. The $\vec{s}$, $A\vec{d}$ and $\vec{s}_{\rm c}$ for each detector are shown in Figure 4. For the jiggle position 0 data, this template subtraction leaves a residual with a standard deviation of $0.6$ MJySr-1, approximately a factor of $3$ better than what was achieved using the standard correction and with no discernible Marchili-like structure. However, this correction works less well for the three jiggle positions that were not present in the Dark Sky observation, which constitute different optical paths through the instrument, and leaves residual Marchili-like structure due to incomplete cancellation. We therefore opt to exclude data taken in jiggle positions $\{1,2,3\}$ from analysis, which increases the statistical uncertainty by a factor of $2$ but removes a likely uncharacteristic systematic error from the measurement.
@@
Here, $\sigma_{I}$ is the per spectral channel uncertainty, and $\delta I_{\rm T}$ is the partial derivative with respect to $T_{\rm rSZe}$. The inverse-squared of this is plotted as a function of spectral channel, and for several different values of $T_{\rm rSZe}$, in Figure 5. We conclude the optimal spectral region is 550–900 GHz, as the sensitivity to the rSZe signal outside of this range is small, and including those data would significantly increase the measurement error. This range also removes the portions of the spectra that may contain significant instrumental systematics.
Figure 5: constraining power on $T_{\rm rSZe}$ estimated by the inverse square of equation 3. The solid black vertical lines highlight where the low- and high-frequency cuts from §III.1 are applied, and outside of which the constraining power is small. While the SZe distortion is spectrally smooth (see Fig. 1), the relativistic corrections exhibit $\sim 100$ GHz wide spectral deviations whose relative amplitudes change over the falling edge of the SZ spectrum. These differential effects provide a lever arm to distinguish gas at different temperatures independently of X-ray information. Hence, the constraining power peaks at  $640$ and  $800$ GHz. This is informative for cutting the low- and high-frequency ends of the FTS bandpass, as the regions with very low constraining power, e.g. $\nu<550$, will contribute disproportionately more noise to the measurement.
III.2 Spectral Noise Estimation
A covariance matrix is estimated using a timewise jackknife resampling that is designed to cancel spatially and spectrally stable signals in our data, e.g., from astrophysical sources or instrument thermal emission, while retaining variations due to statistical noise and properties of the instrument varying on the scan time scale. To generate this estimate, two hundred of the four hundred spectra for a given pointing are multiplied by negative 1, and the mean is computed with:
@@ III.2 Spectral Noise Estimation
Here, $\vec{s}^{*}_{k\pm}$ denotes the sum over all $400$ spectra for a given pointing, where indices with a minus sign are multiplied by negative one and indices with a plus sign are kept the same. The resulting covariance matrix is estimated by computing the outer product of $\vec{s}_{l}$ with itself,
@@
where $\rm T$} denotes the matrix transpose. This process is repeated $2\times 10^{6}$ times (chosen to estimate the covariance matrices to five significant figures) and averaged. We perform these jackknives independently for each detector, jiggle position (only jiggle position 0 is used in the analysis), and observation, producing $28$ unique covariance matrices. Since we expect multiple observations with the same detector to have consistent systematics, we average these by detector down to a total of seven. A representative example showing the correlation matrix, which is estimated from the noise covariance matrix, is shown in Figure 6. The characteristic correlation length implied by the noise covariance is approximately 50 GHz, comparable to the nominal LR spectral resolving power of 25 GHz.
@@
III.3 Unwanted Astrophysical Emission